Show me step by step on how to solve this problem in Excel, thanks.

Tempura, Inc., is considering two projects. Project A requires an investment of $50,000. Estimated annual receipts for 20 years are $20,000; estimated annual costs are $12,500. An alternative project, B, requires an investment of $75,000, has annual receipts for 20 years of $28,000, and has annual costs of $18,000. Assume both projects have zero salvage value and that MARR is 12 percent/year.

a. What is the annual worth of each project?

b. Which project should be recommended?
Answer #1

(a) For each project, Net annual benefit = Annual receipts - Annual cost

Project A: $(20,000 - 12,500) = $7,500

Project B: $(28,000 - 18,000) = $10,000

(i) Cash flow dagram (All values in $)

20000 20000 20000 .. a 3 ... ... 20 Soooo 2500 12500 11500...... 1230D 28000 28000 28000....... 75000 1000 18000 18000 .....

(ii) Present worth (PW) is computed as follows.

PW, Project A ($) = - 50,000 + 7,500 x P/A(12%, 20) = - 50,000 + 7,500 x 7.4694** = - 50,000 + 56,021 = 6,021

PW, Project B ($) = - 75,000 + 10,000 x P/A(12%, 20) = - 75,000 + 10,000 x 7.4694** = - 75,000 + 74,694 = - 306

(b) Since Project A has a positive PW, this should be recommended.

**From PP/A factor table